# Benchmarking LLMs for Political Science: A United Nations Perspective

### v1

@@ 5.4 Task 2: Representatives Voting Simulation
Focused on simulating country-specific voting behavior, this task tests models’ ability to infer issue-specific voting patterns by analyzing how nations prioritize resolution content and contextual geopolitical dynamics. Unlike Task 1, which evaluates proactive alliance-building, Task 2 emphasizes reactive decision-making based on the interplay of national interests, ideological alignment, and external pressures. GPT-4o achieves the highest performance (0.823 Bal. ACC, 0.696 PR AUC), demonstrating a strong ability to model nuanced trade-offs. DeepSeek-V3 (0.724 Bal. ACC, 0.655 PR AUC) and Llama-3.2-3B (0.597 Bal. ACC) show moderate success but struggle with ambiguous cases, while Mistral-7B (0.426 Bal. ACC) performs poorly, reflecting its inability to systematically weigh competing factors. The results highlight LLMs’ potential to simulate diplomatic behavior but reveal significant variance in their capacity to reason about issue-specific voting dynamics.
@@ 5.5 Task 3: Draft Adoption Prediction
Whereas Task 2 centers on individual votes, Task 3 measures document-level outcome prediction, requiring holistic reasoning about all 15 Council members to predict whether a resolution is eventually [Adopted] or [NotAdopted]. GPT-4o shows the best Bal. ACC ($0.677$) and a competitive macro-F1 ($0.363$), while Llama-3.2-3B surpasses others in macro-F1 ($0.402$) but has lower Bal. ACC ($0.597$). The divergence between Bal. ACC and F1 metrics reveals the challenge of modeling adoption mechanics, where understanding Council-wide political dynamics, potential veto threats, and support coalitions play roles.
@@ 5.6 Task 4: Representatives Statement Generation
6 Conclusions
@@
This paper introduces UNBench, the first comprehensive benchmark for evaluating LLMs’ capabilities in political science through UN Security Council records (1994-2024). By designing four interconnected tasks spanning the complete UN resolution lifecycle—co-penholder judgement, representatives voting simulation, draft adoption prediction, and representative statement generation—we provide a more authentic framework for assessing LLMs’ understanding of complex diplomatic dynamics. Our work not only addresses the current gap in LLM evaluation frameworks but also establishes a foundation for future research at the intersection of artificial intelligence and international relations, demonstrating how LLMs could potentially assist in analyzing global governance processes. More applications and detailed analyses can be found in Appendix Benchmarking LLMs for Political Science: A United Nations Perspective.
7 Limitations
Despite UNBench’s contributions, several limitations should be noted. Our dataset is restricted to UN Security Council records from 1994-2024, which may not fully represent broader international relations dynamics. The benchmark currently focuses on English-language documents, potentially missing nuances in diplomatic communications in other UN languages. Additionally, as political landscapes evolve rapidly, historical patterns in our training data may not accurately reflect current diplomatic dynamics. Finally, there exists potential data contamination since LLMs may have been pre-trained on publicly available UN documents. The original UN Security Council records are publicly available, and their interpretation and official authority remain with the United Nations. In this paper, generative AI tools (ChatGPT, Grammarly) are used to fix grammar bugs and typos. As a benchmark for political science, we do not foresee any potential risks that need to be disclosed.
@@ A.1 Subjects Analysis
A.1.1 Author-subject Co-occurrence Analysis
@@ A.2.1 Voting Frequency
A.2.2 Country Relationships revealed within UN Resolutions
Table 3 presents the top 10 pairs of authors and countries that most frequently voted "No" or abstained on draft resolutions. The data reveals that the United States has the highest number of rejections, with Russia rejecting 68 times and China 52 times. This indicates a consistent divide between the U.S. and these two powers, likely reflecting ongoing geopolitical tensions. Similarly, the United Kingdom and France also show high rejection rates from both Russia and China, suggesting a shared stance among Western powers in opposition to certain resolutions proposed by these countries. On the other hand, countries like Germany, Japan, and Italy appear less frequently in the table, with rejections ranging from 13 to 22 times. These smaller states seem to align more often with the major powers but still demonstrate some differences, particularly with China and Russia. Overall, the table highlights significant diplomatic rifts between the Western powers and Russia/China, with frequent rejections indicating key areas of contention in international relations.
@@ A.2.2 Country Relationships revealed within UN Resolutions
Table 4 and Table 5 provide insights into the voting patterns of country pairs within the United Nations, highlighting both strong collaboration and significant divergence in voting behaviors. Table 4 shows that certain country pairs, such as France and the United Kingdom (1,153 joint "Yes" votes), United Kingdom and the United States (1,147 joint "Yes" votes), and France and the United States (1,142 joint "Yes" votes), consistently align on many resolutions, reflecting their close diplomatic and strategic cooperation. In contrast, China and the Western powers also exhibit frequent collaboration, with the China-United Kingdom, China-France, and China-United States pairs each voting together over 1,000 times, indicating areas of common interest despite occasional political differences. On the other hand, Table 5 reveals country pairs that most often did not vote "Yes" together. The China-Russia pair stands out with 69 instances of disagreement, indicating that the two countries share some geopolitical interests. Other pairs, such as Algeria-China (8 times) and Russia-South Africa (6 times), also display divergent voting patterns, reflecting how national and regional interests can influence voting behavior at the UN. These tables underscore the complex dynamics of international diplomacy, where countries may cooperate on certain issues while diverging on others based on their specific interests and priorities.
@@
A.3 Dataset Construction
The overarching goal of this work is to provide a unified and extensive collection of the decision-making process, thereby enabling evaluation of multiple LLM capabilities in a single benchmark. To achieve this, we collect multi-perspective data from the official website and digital libraryUnited Nations Security Council (2024), which archive draft resolutions, voting records, and meeting minutes. Below, we highlight key challenges and our corresponding strategies in constructing $\mathcal{D}$ in the different stages of our benchmark construction.
Data Collection.
Data Conversion.
UN documents are primarily stored in PDF format, making direct ingestion by current LLMs infeasible. We therefore extract and convert the content into plain text. Early attempts using generic Python PDF libraries yielded mixed accuracy due to the unstructured, domain-specific nature of political documents. We applied a LLM-based parser (LlamaParseLlamaIndex (2024)) to handle complex formatting (e.g., multi-column layouts, footnotes, multilingual text).
@@ A.3 Dataset Construction
Data Processing.
(1) Labeling Adopted vs. Unadopted Drafts. Some drafts never become official resolutions (i.e., “unadopted”), lacking a formal resolution ID. We thus inspect the official notes in each draft’s record and cross-verify with the final resolution index to categorize them correctly. (2) Country Name Normalization. Different records refer to the same country with variations (e.g., “United Kingdom” vs. “Kingdom”). We employ a standardized dictionary to unify references to the same country entity across all entries. (3) Metadata Alignment. For each draft $r_{i}$, we compile relevant information—author/sponsor countries, date, issue category, voting breakdown, and meeting transcripts—into a structured format compatible with modern NLP frameworks.

### v2

@@ 5.4 Task 2: Representatives Voting Simulation
Focused on simulating country-specific voting behavior, this task tests models’ ability to infer issue-specific voting patterns by analyzing how nations prioritize resolution content and contextual geopolitical dynamics. Unlike Task 1, which evaluates proactive alliance-building, Task 2 emphasizes reactive decision-making based on the interplay of national interests, ideological alignment, and external pressures. GPT-4o achieves the highest performance (0.823 Bal. ACC, 0.696 PR AUC), demonstrating a strong ability to model nuanced trade-offs. DeepSeek-V3 (0.724 Bal. ACC, 0.655 PR AUC) and Llama-3.2-3B (0.597 Bal. ACC) show moderate success but struggle with ambiguous cases, while Mistral-7B (0.426 Bal. ACC) performs poorly, reflecting its inability to systematically weigh competing factors. The results highlight LLMs’ potential to simulate diplomatic behavior but reveal significant variance in their capacity to reason about issue-specific voting dynamics. We also assess whether model performance varies with respect to recency, potential geopolitical disparities, and different contextual metadata. Detailed results are provided at Appendix C.1, C.2, and C.3.
@@ 5.5 Task 3: Draft Adoption Prediction
Whereas Task 2 centers on individual votes, Task 3 measures document-level outcome prediction, requiring holistic reasoning about all 15 Council members to predict whether a resolution is eventually [Adopted] or [NotAdopted]. GPT-4o shows the best Bal. ACC ($0.677$) and a competitive macro-F1 ($0.363$), while Llama-3.2-3B surpasses others in macro-F1 ($0.402$) but has lower Bal. ACC ($0.597$). The divergence between Bal. ACC and F1 metrics reveals the challenge of modeling adoption mechanics, where understanding Council-wide political dynamics, potential veto threats, and support coalitions play roles. In addition, we analyze GPT-4o’s robustness across time, regional authorship, and metadata conditions for this task. The corresponding evaluations are detailed at Appendix C.1, C.2, and C.3.
@@ 5.6 Task 4: Representatives Statement Generation
6 Potential Applications
The UNBench offers significant value to both LLM researchers and stakeholders in international governance, enabling practical applications and advancing research in geopolitical AI. Below, we outline potential use cases for different stakeholders.
For LLM Researchers:
UNBench provides a rich testbed for advancing research in LLMs, particularly in the context of geopolitical reasoning and time-series analysis: (1) Geopolitical Reasoning: The tasks in the benchmark span a wide range of capabilities, from alliance identification (Task 1) to issue-specific voting prediction (Task 2), offering researchers a comprehensive framework for evaluating and improving LLMs’ understanding of international relations. (2) Temporal Analysis: With data spanning 30 years, UNBench enables time-series tasks such as predicting trends in diplomatic behavior, forecasting shifts in international alliances, or analyzing the impact of historical events (e.g., the end of the Cold War) on UNSC dynamics. For instance, researchers could use the dataset to predict how emerging global issues (e.g., climate change) will influence future resolutions. (3) Fine-Grained Prediction: The benchmark’s focus on multi-choice and generative tasks challenges researchers to develop models that balance precision and creativity. For example, improving ROUGE scores in Task 4 could lead to breakthroughs in generating protocol-compliant diplomatic text. (4) Bias Analysis: UNBench provides an opportunity to study biases in LLMs’ geopolitical reasoning, ensuring that models do not perpetuate stereotypes or oversimplify complex international dynamics.
For UN Stakeholders:
The ability to predict and analyze UNSC decision-making using LLMs also has implications for diplomats, policymakers, and international organizations: (1) Draft Adoption Forecasting: By predicting whether a draft resolution will be adopted (Task 3), stakeholders can proactively adjust negotiation strategies, allocate resources more effectively, and build coalitions to maximize the likelihood of success. For example, knowing that a climate resolution is unlikely to pass could prompt earlier lobbying efforts or revisions to the draft. (2) Voting Behavior Simulation: Simulating country-specific voting behavior (Task 2) allows stakeholders to anticipate the positions of key nations, identify potential allies or opponents, and tailor diplomatic outreach accordingly. This could be particularly useful for smaller nations or NGOs seeking to navigate complex geopolitical landscapes.
7 Conclusion and Future Work
@@
This paper introduces UNBench, the first comprehensive benchmark for evaluating LLMs’ capabilities in political science through UN Security Council records (1994-2024). By designing four interconnected tasks spanning the complete UN resolution lifecycle, we provide a more authentic framework for assessing LLMs’ understanding of complex diplomatic dynamics. It not only addresses the current gap in LLM evaluation frameworks but also establishes a foundation for future research between artificial intelligence and international relations, demonstrating how LLMs could potentially assist in analyzing global governance processes. More detailed analyses can be found at Appendix in the extended version (liang2025benchmarking).
Future work can extend UNBench to multilingual settings for evaluating models across diverse diplomatic languages. Researchers may also use UNBench to study more realistic multi-agent simulations where LLMs act as countries, as well as multimodal inputs that incorporate maps, timelines, or historical context. Another direction is Constructing richer country profiles to enhance grounding and strategic consistency when models simulate state behavior.
Acknowledgments
This material is based upon work supported by NSF awards (SaTC-2241068, IIS-2506643, and POSE-2346158), a Cisco Research Award, and a Microsoft Accelerate Foundation Models Research Award. The views and conclusions contained in this document are those of the authors and should not be interpreted as necessarily representing the official policies, either expressed or implied, of the National Science Foundation.
@@ A.1 Subjects Analysis
Author-subject Co-occurrence Analysis
@@ A.2 National-Wised Analysis
Country Relationships revealed within UN Resolutions
Table 3 presents the top 10 pairs of authors and countries that most frequently voted ”No” or abstained on draft resolutions. The data reveals that the United States has the highest number of rejections, with Russia rejecting 68 times and China 52 times. This indicates a consistent divide between the U.S. and these two powers, likely reflecting ongoing geopolitical tensions. Similarly, the United Kingdom and France also show high rejection rates from both Russia and China, suggesting a shared stance among Western powers in opposition to certain resolutions proposed by these countries. On the other hand, countries like Germany, Japan, and Italy appear less frequently in the table, with rejections ranging from 13 to 22 times. These smaller states seem to align more often with the major powers but still demonstrate some differences, particularly with China and Russia. Overall, the table highlights significant diplomatic rifts between the Western powers and Russia/China, with frequent rejections indicating key areas of contention in international relations.
@@
Table 4 and Table 5 provide insights into the voting patterns of country pairs within the United Nations, highlighting both strong collaboration and significant divergence in voting behaviors. Table 4 shows that certain country pairs, such as France and the United Kingdom (1,153 joint ”Yes” votes), United Kingdom and the United States (1,147 joint ”Yes” votes), and France and the United States (1,142 joint ”Yes” votes), consistently align on many resolutions, reflecting their close diplomatic and strategic cooperation. In contrast, China and the Western powers also exhibit frequent collaboration, with the China-United Kingdom, China-France, and China-United States pairs each voting together over 1,000 times, indicating areas of common interest despite occasional political differences. On the other hand, Table 5 reveals country pairs that most often did not vote ”Yes” together. The China-Russia pair stands out with 69 instances of disagreement, indicating that the two countries share some geopolitical interests. Other pairs, such as Algeria-China (8 times) and Russia-South Africa (6 times), also display divergent voting patterns, reflecting how national and regional interests can influence voting behavior at the UN. These tables underscore the complex dynamics of international diplomacy, where countries may cooperate on certain issues while diverging on others based on their specific interests and priorities.
@@ Appendix B Detailed Results of UNBench
Appendix C Additional Analyses
C.1 Temporal Robustness Evaluation
To assess the temporal generalization of language models, we divide the test set based on GPT-4o’s training cutoff (October 2023) into pre-2023 (seen) and post-2023 (unseen) subsets. This helps us evaluate whether models perform worse when facing emerging, unseen geopolitical developments.
Task 2: Voting Simulation. GPT-4o shows strong performance overall, but drops notably on unseen (post-2023) data. Balanced accuracy falls from 0.675 on seen examples to 0.528 on unseen ones, highlighting the challenge of temporal generalization.
Model Accuracy Bal. ACC AUC F1
GPT-4o 0.922 0.677 0.731 0.363
GPT-4o-seen 0.924 0.675 0.727 0.360
GPT-4o-unseen 0.907 0.528 0.640 0.376
Task 3: Draft Adoption Prediction. In contrast, GPT-4o maintains high accuracy across time splits, suggesting the model’s adoption reasoning is less sensitive to temporal shifts.
Model Accuracy Bal. ACC AUC F1
GPT-4o 0.968 0.823 0.823 0.686
GPT-4o-seen 0.976 0.816 0.816 0.737
GPT-4o-unseen 0.962 0.828 0.828 0.656
C.2 Geopolitical Group Analysis
We further investigate whether model behavior differs depending on the geopolitical affiliation of draft resolution authors, specifically contrasting Western European and Others Group (WEOG) members with non-WEOG states. This evaluation provides a fairness-oriented lens on LLM behavior.
Task 2: Voting Simulation. GPT-4o performs better on resolutions drafted by non-WEOG countries (balanced accuracy: 0.690) than on those authored by WEOG states (0.537). The discrepancy suggests that the model may struggle to simulate diplomatic reactions to Western-sponsored resolutions.
Group Bal. ACC Precision Recall PR AUC
WEOG 0.537 0.441 0.358 0.524
Non-WEOG 0.690 0.384 0.690 0.346
Task 3: Draft Adoption Prediction. Results are more stable across regions for adoption outcomes, with GPT-4o achieving comparable balanced accuracy on WEOG (0.815) and non-WEOG (0.824) drafts.
Group Bal. ACC Precision Recall PR AUC
WEOG 0.815 0.778 0.636 0.714
Non-WEOG 0.824 0.700 0.667 0.693
C.3 Contextual Input Variation
We test whether including metadata like dates or authorship improves LLM decision-making. This is especially relevant for interpreting real-world diplomatic context.
Task 2: Voting Simulation. Adding author metadata improves balanced accuracy (from 0.677 to 0.684) and G-Mean (from 0.729 to 0.753), suggesting model predictions benefit from geopolitical signaling embedded in authorship.
Setting Bal. ACC Recall G-Mean
GPT-4o 0.677 0.677 0.729
GPT-4o-dates 0.673 0.673 0.719
GPT-4o-authors 0.684 0.684 0.753
Task 3: Draft Adoption Prediction. Adding date or author metadata boosts both recall and G-Mean, with the strongest performance under the date-enhanced setting.
Setting Bal. ACC Recall G-Mean
GPT-4o 0.823 0.660 0.807
GPT-4o-dates 0.888 0.811 0.885
GPT-4o-authors 0.859 0.769 0.854
C.4 Model Scaling Analysis
Task 1 Task 2 Task 3 Task 4
Model (1/2) (1/5) Bal. ACC PR AUC Bal. ACC Mac. F1 ROUGE Cosine Sim.
Llama-3.2-1B 0.581 0.269 0.546 0.185 0.320 0.326 0.033 0.329
Llama-3.2-3B 0.578 0.297 0.597 0.385 0.597 0.402 0.041 0.290
Qwen2.5-7B 0.642 0.293 0.699 0.375 0.578 0.241 0.201 0.623
Mistral-7B 0.563 0.281 0.426 0.268 0.529 0.140 0.194 0.575
Llama-3.1-8B 0.665 0.379 0.530 0.168 0.357 0.359 0.039 0.355
DeepSeek-V3-70B 0.695 0.422 0.724 0.655 0.668 0.351 0.207 0.623
GPT-4o ( 200B) 0.726 0.464 0.823 0.696 0.677 0.363 0.199 0.619
Table 10: Model scaling comparison across four tasks. Metrics match those in Table 2. Best results are in bold, second-best are underlined.
We compare performance across models of increasing parameter scale to assess how size correlates with capability across tasks. While model size generally aligns with stronger performance, architectural design and training data also have noticeable effects.
Larger models, particularly GPT-4o and DeepSeek-V3, show consistent strength across tasks, with notable gains in political reasoning, coherence, and generation fidelity. However, smaller models like Qwen2.5-7B occasionally match or outperform larger models (e.g., in embedding-based similarity), indicating that scale alone does not guarantee task superiority.
C.5 Evaluation Strategy Enhancements (Task 4)
To supplement automatic metrics (e.g., ROUGE, cosine similarity), we can incorporate enhanced qualitative assessments like:
  • LLM-as-a-Judge: GPT-4 evaluates outputs on tone, coherence, justification quality, and alignment with national interests.
  • Expert Review: Human evaluators from political science backgrounds will provide qualitative ratings and thematic annotations.
These strategies offer richer insights beyond n-gram or embedding-based overlap, and will be used to further benchmark models on Task 4 in future iterations of UNBench.